# SKYSURF IX - The Cosmic Optical and Infrared Background from Integrated Galaxy Light Measurements

### v1

@@ 6.3 Prospects For Future Studies of the EBL
Outside of the optical and NIR regimes, next generation facilities such as the square kilometer array (SKA) have the potential to unlock new wavelength regimes to precisely study the EBL. The CRB still suffers from the disagreement between upper limits established by direct detection techniques (Fixsen et al., 2011) and the integration of source counts (Gervasi et al., 2008; Vernstrom et al., 2011; Tompkins et al., 2023). The SKA has the potential to rule out significant populations of undetected faint or clustered sources, which had been proposed as a cause of previous disagreements in studies of the COB, as a cause for the current status of radio EBL measurements.
@@ 7 Summary
In this work we present the conclusions of the SKYSURF HST legacy archival program in which we produce the most precise measurements of the IGL to date. We produced over 16,000 mosaics starting with calibration level 3 image FLC and FLT image products. We have processed a significant amount of HST data using the same processing pipeline to produce consistent mosaics and object catalogs. Across 16.8 of area we reach completeness limits of AB 24-28.5 mag in different exposure time ranges. By rigorously filtering for problematic fields in biased HST data, we are able to produce reliable galaxy number counts across 17 of 22 HST filters from which mosaics were constructed. We use our HST number counts alongside WAVES and DEVILS to measure the IGL at 17 different wavelengths between 0.3-1.6 m by magnitude zero-point correcting the WAVES and DEVILS number counts to the HST filter set. We then extensively discuss the significant sources of error affecting the IGL, Poisson, cosmic variance, and magnitude zero point uncertainties. We then comment on the self-consistency of our HST IGL measurements in adjacent HST filters.
@@ 8 Acknowledgments
This work was made possible by the following grants, HST programs AR-09955 and AR-15810 provided by NASA through the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Timothy Carleton is thankful for support from the Beus Center for Cosmic Foundations.

### v2

@@ 6.3 Prospects For Future Studies of the EBL
Outside of the optical and NIR regimes, next generation facilities such as the square kilometre array (SKA) have the potential to unlock new wavelength regimes to precisely study the EBL. The CRB still suffers from the disagreement between upper limits established by direct detection techniques (Fixsen et al., 2011) and the integration of source counts (Gervasi et al., 2008; Vernstrom et al., 2011; Tompkins et al., 2023). The SKA has the potential to rule out significant populations of undetected faint or clustered sources, which had been proposed as a cause of previous disagreements in studies of the COB, as a cause for the current status of radio EBL measurements.
The COB encodes the CSFH and using the semi-analytical modelling software ProSpect Robotham et al. (2020). In Figure 23 we show that using a prescribed cosmic star formation history we are able predict the COB. In combination with the CSFH, producing an accurate model requires some understanding of ongoing evolutionary processes within galaxies such as a stellar initial mass function (IMF), evolving metallicity history, and canonical model for dust attenuation. We plan to use our COB measurements in combination with the CRB measurements of Tompkins et al. (2023) to reverse this process and predict the CSFH in future work. As such, we only provide a brief discussion of EBL modelling and the CSFH. We adopt the BC03 stellar libraries Bruzual and Charlot (2003) and dust attenuation templates of Dale et al. (2014). We also maintain a constant Chabrier (2003) IMF throughout. In this demonstration, our closed box metallicity evolution begins at a metallicity value of and reaches solar metallicity at z=0. To generate mock SEDs, we keep most ProSpectSED function parameters at default values. For a detailed discussion of all parameters see the relevant documentation. We produce EBL predictions using these inputs via the summation of the appropriately redshifted ProSpect SEDs starting with a look-back time of 13.5 GYR and ending at a look-back time of GYR.
@@ 7 Summary
In this work we present the conclusions of the SKYSURF HST legacy archival program in which we produce the most precise measurements of the IGL to date. We produced over 16,000 mosaics starting with calibration level 3 image FLC and FLT image products. We have processed a significant amount of HST data using the same processing pipeline to produce consistent mosaics and object catalogs. Across 16.8 of area we reach completeness limits of AB 24-28.5 mag in different exposure time ranges. By rigorously filtering for problematic fields in biased HST data, we are able to produce reliable galaxy number counts across 17 of 22 HST filters from which mosaics were constructed. We use our HST number counts alongside WAVES and DEVILS to measure the IGL at 17 different wavelengths between 0.3-1.6 m by magnitude zero-point correcting the WAVES and DEVILS number counts to the HST filter set. We extensively discuss the significant sources of error affecting the IGL, Poisson, cosmic variance, and magnitude zero point uncertainties. We test our sample for systematic biases which could arise as a function of either galactic or ecliptic latitude. We then comment on the self-consistency of our HST IGL measurements in adjacent HST filters.
@@ 8 Acknowledgments
This work was made possible by the following grants, HST programs AR-09955 and AR-15810 provided by NASA through the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Timothy Carleton is thankful for support from the Beus Center for Cosmic Foundations. Finally, we thank our reviewer whose comments and suggestions strengthened the presentation and robustness of our results.
@@ 9 Software
10 Data Availability
In accordance with the SKYSURF grant conditions we make our source catalogs publicly available upon request to the corresponding author. The catalogs contain the entire output of the ProFound source detection code, for a description of ProFound columns see the ProFound documentation. Here we describe the added columns.
Starflag - Flag specifying if an object is a identified as a star. A value of 0 specifies an object as a galaxy and a value of 1 as a star.
Secondary Flag - A flag specifying whether the object is used in the production of counts. A value of 0 is a good object identified as a galaxy, while 1 is given to stars, 2 to suspected artifacts, 4 to given to objects falling within exclusion radii around stars such as diffraction spikes, and a value of 3 is given to objects brighter than AB 15 mag which are not used and have their surroundings masked.
Exptime - The stacked exposure time at the centre of each object.
wht - The weightmap value, or number of good pixels in the stack, at the centre of each object. Objects with a value of zero are possible noise and not included.
Region - An integer, identifying which exposure time bin, eg, 200-400 (1), 400-800 (2), an object falls into, used for keeping track of areas.
Area - The total unmasked area of each mosaic.
sub area - The total area of unmasked pixels in the associated exposure time range.
fraction - The fraction of area covered by the corresponding exposure time range.
visit - ID flag given to specify which mosaic an object comes from.
corr mag - Galactic extinction corrected magnitudes.